what is the measure of each exterior angle of a regular 15 sided polygon?
Answer by Don2xmalabag(6)

exterior angle = 360/n where n is the number of sides
Ext. Angle = 360/15
Ext.Angle = 24
*so the measure of the exterior angle of a 15-sided polygon is 24 degrees.